One number is 12 more than another. The sum of the smaller number and twice the larger number is 39. Find the larger number.
Answer by scott8148(6628)

let x=larger, so x-12=smaller

"The sum of the smaller number and twice the larger number is 39" ___ x-12+2x=39

3x-12=39